# When in Doubt, Cascade: Towards Building Efficient and Capable Guardrails

### v1

@@ 7.1 Limitations
We acknowledge that our taxonomy (even the full version in Appendix B) cannot cover all groups which are salient for social bias. This taxonomy is a point-in-time artifact, and we note that social bias dynamically evolves over time. Because social harms are the product of context-dependent classification systems with deep historical roots and are socially and morally charged, careful attention must be paid to the choices made (such as the groups included in a taxonomy) during development of the detectors [1]. Concretely, we understand that generated counter-speech may not always be reflective of real-world harm. One way to combat this, though expensive, is via human annotations with true counter-speakers [26]. Another direction is to leverage LLMs themselves to provide feedback on a given counter-speech example [20]. We are experimenting with both avenues in the near future.
@@ 7.2 Future Work
In order to combat overconfidence, we are considering conformal prediction approaches [40]. These approaches quantify uncertainty in a model’s prediction by constructing predictive sets (as opposed to singleton labels, in the case of our detectors) with guaranteed frequentist coverage probabilities. Specifically, we are looking into the regularized adaptive prediction sets approach [34, 2] which, in addition to providing coverage guarantees, produces larger (or non-singleton in the case of our detectors) prediction sets for difficult instances and smaller (singleton) sets for easier to classify examples. Then, on the instances for which our models produce these larger sets, we are exploring is the idea of a “confident-cascade" wherein we adopt a collaborative method and offload these inputs to Llama-Guard. We conjecture this may provide a reasonable trade-off between computational efficiency and performance.
@@
Currently, detectors are being trained in a supervised fine-tuning method with the standard binary cross-entropy loss. Given that our synthetic data is generated in a contrastive manner, we will consider training with contrastive loss [21]. Additionally, we are looking into directly training a multi-head detector, which eliminates the need for two models in our cascade approach.
@@ Appendix A Social Bias Detector
As mentioned in Section 1, we trained a social-bias-detector by fine-tuning BERT [10]. In particular, we took the uncased BERT model from HuggingFace111111https://huggingface.co/google-bert/bert-base-uncased. During training, we use a batch size of 16, we start with a learning rate of , and we train for 50 epochs, taking the best model with respect to validation f1 score. For reference, the total training time was a few hours on a single A100 GPU.
@@ A.1 Training Data
We started with the Latent Hatred [12] dataset, which is a benchmark that was specifically designed for implicit hate speech. Then, we use the 20 NewsGroups dataset [24] in order to add some out of distribution data. Note that we deliberately use this dataset in the hopes of increasing the proportion of negative (i.e., benign) labeled data in our training set. Third, we use a dataset from a work titled Identifying Implicitly Abusive Remarks about Identity Groups using a Linguistically Informed Approach [43], as this dataset attempted to decouple harmful intent from mention of specific group identities. In fact, this concept is related to the use-mention distinction that we discuss throughout the paper. We refer to this dataset as “Identity Groups" in Table 5 below. Finally, we add in a subset of the CivilComments dataset [8], taking only samples which have an identity_attack column value of greater than 0.5, which we believe corresponded to implicitly hateful comments.

### v2

@@ 7.1 Limitations
We acknowledge that our taxonomy (even the full version in Appendix B) cannot cover all groups which are salient for social bias. This taxonomy is a point-in-time artifact, and we note that social bias dynamically evolves over time. Because social harms are the product of context-dependent classification systems with deep historical roots and are socially and morally charged, careful attention must be paid to the choices made (such as the groups included in a taxonomy) during development of the detectors (Achintalwar et al. 2024). Concretely, we understand that generated counter-speech may not always be reflective of real-world harm. One way to combat this, though expensive, is via human annotations with true counter-speakers (Mun et al. 2024). Another direction is to leverage LLMs themselves to provide feedback on a given counter-speech example (Jones et al. 2024). We are experimenting with both avenues in the near future.
@@ 7.2 Future Work
In order to combat overconfidence, we are considering conformal prediction approaches (Vovk, Gammerman, and Saunders 1999). These approaches quantify uncertainty in a model’s prediction by constructing predictive sets (as opposed to singleton labels, in the case of our detectors) with guaranteed frequentist coverage probabilities. Specifically, we are looking into the regularized adaptive prediction sets approach (Romano, Sesia, and Candes 2020; Angelopoulos et al. 2021) which, in addition to providing coverage guarantees, produces larger (or non-singleton in the case of our detectors) prediction sets for difficult instances and smaller (singleton) sets for easier to classify examples. Then, on the instances for which our models produce these larger sets, we are exploring is the idea of a “confident-cascade” wherein we adopt a collaborative method and offload these inputs to Llama-Guard. We conjecture this may provide a reasonable trade-off between computational efficiency and performance.
@@
Currently, detectors are being trained in a supervised fine-tuning method with the standard binary cross-entropy loss. Given that our synthetic data is generated in a contrastive manner, we will consider training with contrastive loss (Khosla et al. 2021). Additionally, we are looking into directly training a multi-head detector, which eliminates the need for two models in our cascade approach.
@@ Appendix A Social Bias Detector
As mentioned in Section 1, we trained a social-bias-detector by fine-tuning BERT (Devlin et al. 2019). In particular, we took the uncased BERT model from HuggingFace111111https://huggingface.co/google-bert/bert-base-uncased. During training, we use a batch size of 16, we start with a learning rate of , and we train for 50 epochs, taking the best model with respect to validation f1 score. For reference, the total training time was a few hours on a single A100 GPU.
@@ A.1 Training Data
We started with the Latent Hatred (ElSherief et al. 2021) dataset, which is a benchmark that was specifically designed for implicit hate speech. Then, we use the 20 NewsGroups dataset (Mitchell 1999) in order to add some out of distribution data. Note that we deliberately use this dataset in the hopes of increasing the proportion of negative (i.e., benign) labeled data in our training set. Third, we use a dataset from a work titled Identifying Implicitly Abusive Remarks about Identity Groups using a Linguistically Informed Approach (Wiegand, Eder, and Ruppenhofer 2022), as this dataset attempted to decouple harmful intent from mention of specific group identities. In fact, this concept is related to the use-mention distinction that we discuss throughout the paper. We refer to this dataset as “Identity Groups” in Table 5 below. Finally, we add in a subset of the CivilComments dataset (Borkan et al. 2019), taking only samples which have an identity_attack column value of greater than 0.5, which we believe corresponded to implicitly hateful comments.